RTR Inc.'s sales last year were $2,750,000 (all on credit), and its net profit margin was 12%. Its inventory turnover was 5.0 times during the year, and its days sales outstanding was 32 days. Its annual cost of goods sold was $2,200,000. The firm had fixed assets totaling $800,000. RTR's payables deferral period is 45 days.

Assuming RTR holds negligible amounts of cash and marketable securities (zero), calculate

A.) Total assets turnover

B.) Return on assets
Answer #1

A). Inventory Turnover Ratio = Sales / Inventory

5 = $2,750,000 / Inventory

Inventory = $2,750,000 / 5 = $550,000

DSO = 365 / [Sales / Accounts Receivable]

32 = 365 / [$2,750,000 / Accounts Receivable]

$2,750,000 / Accounts Receivable = 365 / 32

Accounts Receivable = $2,750,000 / 11.40625 = $241,095.89

Total Assets = Current Assets + Fixed Assets

= $550,000 + $241,095.89 + $800,000 = $1,591,095.89

Total Assets Turnover = Sales / Total Assets = $2,750,000 / $1,591,095.89 = 1.73 times

B). Return on Assets = Net Income * Total Asset Turnover

= 0.12 * 1.73 = 0.2074, or 20.74%